For DPU of 0.02120, Determine Z equiv , Z STand estimate the ppm defect rate

  1. For DPU of 0.02120, Determine Z equiv , Z STand estimate the ppm defect rate
Answer #1

here as proportion of defective=0.0212 ; therefore z score will lie between 0.0212/2=0.0106 and (1-0.0212/2=0.9894) fraction

for which z equivalent =2.30

estimated ppm defect rate =0.0212*103 =21.2 ppm defects